OpenStudy (anonymous):

Probability: Please Help :( a bag holds 3 colored balls, 1 of which is a bright purple ball. Ramona picks one and puts it back in the bag. Carlos picks one. What is the probability, that they BOTH will pick a purple ball?

OpenStudy (anonymous):

they both have a 1/3 chance of picking the purple ball so p1p2=1/3(1/3)=1/6 chance

OpenStudy (anonymous):

I thought the same thing. Either 1/3 or 1/6 for both of them. But my answer doesn't work, so its wrong :/

OpenStudy (anonymous):

Okay so the chance to pull out a purple ball is clearly \(1/3\). Now, since Ramona put's the ball back in after picking one, it makes both events independent. So in the end, it's just the probabilities multiplied: \(1/3 \cdot 1/3 = 1/9\)

OpenStudy (anonymous):

It said to multiply but I thought it made each number, making 1/3, 2/6 when reduced would just go back to 1/3. BUT that's what it meant!!! Thank you sooo much! :D

OpenStudy (anonymous):

The only numbers multiplied by itself and resulting in itself is \(0\) and \(1\)